OpenStudy (anonymous):

Another inequality...these bad boys are lethal, need help! Solve the inequality x + 8<5 or -8x<16

OpenStudy (zzr0ck3r):

1st equation subtract 8 from both sides 2nd equation divide both sides by -8 and flip the inequality

OpenStudy (anonymous):

ok and what would the interval notation look like

OpenStudy (zzr0ck3r):

(-infnity,-3)U(-2,infinity)

OpenStudy (anonymous):

hmm i was sure it was -3,-2 but ight thanks for the help:)

OpenStudy (anonymous):

Solve both for \(x\) and *combine.* $$x+8<5\\x+8-8<5-8\\x<-3\\x\in(-\infty,-3)$$and$$-8x<16\\\frac{-8x}{-8}>\frac{16}{-8}\\x>-2\\x\in(-2,\infty)$$hence \(x\in(-\infty,-3)\cap(-2,\infty)=\emptyset\) or, equivalently, \(-2<x<-3\) which no real number satisfies...

OpenStudy (zzr0ck3r):

no this is not right

OpenStudy (zzr0ck3r):

or!

OpenStudy (zzr0ck3r):

or not and

OpenStudy (zzr0ck3r):

(-infnity,-3)U(-2,infinity) not (-infnity,-3)n(-2,infinity)

OpenStudy (zzr0ck3r):

it just means x needs to satisfy one equation not both @oldrin.bataku

OpenStudy (anonymous):

@zzr0ck3r trust me I know that. I misread 'or' as 'and'. It would correctly be \(x\in(-\infty,-3)\cup(-2,\infty)\)

OpenStudy (zzr0ck3r):

right, that's why I told you:) I assume you would know it....